# Feeling Optimistic? Ambiguity Attitudes for Online Decision Making

### v2

@@ Abstract
Due to the complexity of many decision making problems, tree search algorithms often have inadequate information to produce accurate transition models. Robust methods, designed to make safe decisions when faced with these uncertainties, often overlook the impact expressions of uncertainty have on how the decision is made. This work introduces the Ambiguity Attitude Graph Search (AAGS), advocating for more precise representation of ambiguities (uncertainty from a set of plausible models) in decision making. Additionally, AAGS allows users to adjust their ambiguity attitude (or preference), promoting exploration and improving users’ ability to control how an agent should respond when faced with a set of valid alternatives. Simulation in a dynamic sailing environment shows how highly stochastic environments can lead robust methods to fail. Results further demonstrate how adjusting ambiguity attitudes better fulfills objectives while mitigating this failure mode of robust approaches. Because this approach is a generalization of the robust framework, these results further demonstrate how algorithms focused on ambiguity have applicability beyond safety-critical systems.
@@ I Introduction
As autonomous agents navigate more complex environments, they are confronted with increasing model uncertainty. This is especially important with Monte Carlo tree search techniques, such as upper confidence trees (UCT) [1]. Tree search algorithms sample a blackbox simulator to produce transition models for each state encountered. However, many of these will be inaccurate due to limitations on sampling and the tree structure of the graph [2]. Consider a home-service robot, many diverse encounters and long operational periods are required before enough data is present to model a given user. This leaves robots with too little information to reliably predict outcomes—that is to say the uncertainty is non-measurable [3] or ambiguous. Conditioned on some prior information, these ambiguities can be thought of as equally valid models of a system. For example, when asking a robot to get a cup, we may be surprised to find they return with a glass rather than a mug for our hot tea. By not specifying the drink ahead of time, the robot was left with a number of options and no distinguishing information. Failure to consider these ambiguities in the context of decision making can distort the perception of uncertainty and leaves robots oblivious to potential fault conditions.
@@
Some more recent works have begun to consider ambiguity in a limited sense. These approaches generally fit under the auspices of robust decision making. Robust decision making tools construct a set of models and seek to maximize the lower bound on a state’s value [7, 8]. Representations of ambiguity in this context are generally limited to the use of confidence intervals [2, 9] or chance constraints on undesirable outcomes [10]. These representations neglect other information conveyed by the transition model. For example, by considering the distribution of outcomes more explicitly, we can arrive at more informed decisions [11]; coupled with representations of ambiguity we can better assess the knowledge carried by a decision maker. Additionally, the conservatism of robust methods limits their use to safety critical systems. In many other applications (e.g., service robots), giving decision makers the option to trade-off between robust policies and more exploratory or optimistic policies when faced with ambiguity can lead to more desirable outcomes (Fig. 1).
@@ II Ambiguity MDP
Markov decision processes (MDP) have seen broad use formulating decision problems [14] and have a substantial base of literature from which to build tools. The key issue with MDPs is the assumption of one true transition model. For methods which produce approximate solutions, this model is hidden. Multiple transition models, then, may be inferred, implying significant ambiguity when data is limited. This work builds on the concept ambiguous Markov decision processes (AMDP) [15, 16], which frame the problem with a set of models. Define the AMDP as $\mathcal{M}\coloneqq\langle\mathcal{S},\mathcal{A},\mathcal{T},R,\gamma\rangle$. Let $s\in\mathcal{S}$ be a state in the set of states and $a\in\mathcal{A}$ be an action in the action space. The reward $R(s,a,s^{\prime})$ defines the optimization criterion with respect to a specified state transition. The variable $\gamma$ acts a temporal discount factor. Lastly and central to the AMDP formulation, $T(s^{\prime}|s,a)\in\mathcal{T}$ is a model among the set valid of transition models. In this work the set of models is represented as a special class of random sets—referred to as belief functions (different from beliefs in POMDP literature [^1] Beliefs are defined specifically with respect to state uncertainty. Belief functions refer to mass functions whose elements may support more than one outcome.). As opposed to previous AMDP formulations, use of belief functions are better suited to discrete distributions than cloud models as used by Saghafian [15]. Additionally, these belief functions generalize Bayesian mass functions, allocating mass to multiple hypotheses [4]. Thus they permit more precise expressions of ambiguity than those derived from Bayesian models (e.g., the AMDP of Baürle and Rieder [16]).
@@
These bounds on probability are demonstrated in Figure 2.a where each mass is known to ±$\pm\epsilon$. Figure 2.b captures the distribution as a belief function where mass can now support multiple hypotheses.
@@ III Methods
The goal of this work is to motivate the broader study of ambiguity in decision making. The proposed ambiguity attitude graph search algorithm builds on this by better representing ambiguity present in decision making, and generalizing robust decision making to include ambiguity attitudes. As with many sampling based methods [1, 2, 14], agents supply an $(s,a)$-pair to a simulator. This simulator returns observations in the form of a state transition and reward $\omega\coloneqq(s^{\prime},r)$. Using these observations, the agent constructs an estimate of the transition model for each $(s,a)$-pair. However, multiple underlying transition models could have produced these observations. Lacking additional information, there is ambiguity and the model estimate should instead be a set of models. Whereas prior methods, such as graph based optimistic planning (GBOP) [2], assume ambiguity is adequately described by confidence intervals, we contribute a novel way to generate belief functions from these intervals to more precisely represent the distribution of mass. Whereas robust methods assume the lower bound should be used to arbitrate models, AAGS uses ambiguity attitudes to balance robustness and optimism.
@@ III-A Computing Belief Functions
Approximate decision making algorithms generally define their optimality either with respect to the budget or the accuracy of estimated models [9]. In the case of fixed confidence approaches, a confidence interval over some value is specified to accuracy $\epsilon$ given confidence $\delta$. The idea is that a quantity $z$ is approximated with $\hat{z}$ such that $P(|z-\hat{z}|<\epsilon)\geq 1-\delta$. Whereas these bounds are generally applied to the reward or value estimates, AAGS bounds the mass terms instead to evaluate the distribution more directly. Focusing on the entire distribution in this way has been shown to be more expressive and provide a better understanding of the model [11].
When there are limited samples, it is unlikely the constraints for a fixed confidence approach will be satisfied. We therefore develop a “floating confidence” description of the set of transition models for this case. Thus users can understand the confidence with which a distribution of desired accuracy may be interpreted. A belief function is then evaluated using linear optimization. Using the confidence, the set of models inferred from samples is discounted against the set of all models to arrive at a distribution expressing the ambiguity.
Begin by assuming the transition model is represented by a multinomial distribution, known to $(\epsilon,\delta)$. Our approach is agnostic to how these are obtained (for clarity, implementation details are presented in the Appendix). We wish to redistribute the $\epsilon$ mass from each outcome to ambiguous mass terms, generating a set of transition models. Define $\Theta$ as the set of outcomes $(s^{\prime},r)$ from our state transition, and let $2^{\Theta}$ be the power set of all such outcomes. Let each proposition of size one be $\omega_{i}\in\Theta$ and propositions of size two or more be $\theta_{j}\in\{2^{\Theta}\setminus\Theta\}$. The accuracy constrains belief functions such that the lower and upper bounds are, respectively, $Bel(\omega)=p_{\omega}-\epsilon$ and $Pl(\omega)=p_{\omega}+\epsilon$, $\forall\omega\in\Theta$. Because these are probabilities, they are clipped to 0 and 1, as necessary.
@@ III-B Ambiguity Attitude Graph Search
AAGS is a graph based planner, thus using samples more efficiently than tree methods [2]. Assuming the simulator is much more expensive than the algorithm itself, the cost of any added computation are far outweighed by the reduced sample complexity. Cyclic graphs can occur during search and there are no constraints on connectivity.
@@ IV Results
We compare AAGS against two algorithms. First, UCT [1] serves as a baseline. Because UCT does not consider sets of model in its value estimates, it is ambiguity neutral by default. UCT was to tuned to have a rollout horizon of 25 time steps and exploration parameter $c=8$. Next, GBOP [2] was used to compare against robust planning. Both AAGS and GBOP were given a horizon of 50 time steps. All algorithms were allocated a budget of 500 samples.
@@ IV-A Expression of Ambiguity
First, let us draw attention to the parameter $c$ used by UCT. This exploration parameter affects its search, not the representation of value during decision making. Thus, at leaf nodes there is no concept of how uncertain the reward backpropagated to parents is. Robust algorithms have begun to consider this uncertainty, however, they assume the lower bound is most relevant to the decision. This motivates the use of ambiguity attitudes to reflect how willing the agent should be to accept unknown gambles and factor this into the representation of the problem. One impact, is that users can control the extent to which the agent explores its environment. This is demonstrated in Figure 1 where we compare the role of planning horizons with ambiguity attitude for AAGS. We see that in some cases long planning horizons are necessary to for a robust attitude ($\alpha=0$) to identify high value targets. In which case, they move directly towards the target. Given a short planning horizon and increasingly optimistic attitude ($\alpha\rightarrow 1$), however, agents increasingly seek alternatives. They move towards ambiguous regions to learn about their environment. This results in more exploratory behavior despite shorter search horizons.
@@ IV-B Sailing
We first look to the role of ambiguity in avoiding unsafe outcomes. Interestingly GBOP, a robust algorithm struggles with this (Fig. 4.a). We postulate this is because the fixed confidence models employed are relatively uninformative in highly stochastic environments such as this. The UCT and robust ($\alpha=0$) AAGS algorithms, however, almost completely avoid such negative outcomes. Though this number does trend up slightly with increases in initial distance (from the goal), as agents are more likely to start in the boundary. Furthermore, we see that by adjusting the ambiguity attitude of AAGS, we can directly control the willingness of an agent to gamble on low quality outcomes (Fig. 4.b). This is further reflected by the rewards (Fig. 5). For moderate attitudes, an agent can effectively balance reward seeking with chances of hitting the reefs. This can lead to improvements over UCT by $\sim$10%. Conversely, a robust attitude may significantly underperform. Similarly, an optimistic one may gamble too much and degrade outcomes.
Figure 5: (a) Reward for UCT and robust solutions (AAGS at $\alpha=0$ and GBOP) and (b) the reward achieved by AAGS for different values of $\alpha$. Focusing on raising the lower bound, robust attitudes achieve lower reward. Conversely, slightly positive attitudes can outperform ambiguity neutral methods (UCT). When attitudes are too optimistic ($\alpha=1$), they take unnecessary gambles and underperform.
@@ V Conclusions
The increasing complexity of tasks we assign to robots entails ambiguity on the part of the decision maker. This work explores how a decision maker infers multiple models. While existing methods do this in varying capacities, this work aims to demonstrate how more explicit representations of ambiguity contribute to understanding uncertain decisions. In doing so, this lays the groundwork for novel methods to multiple model problems.
Additionally, this work demonstrates how ambiguity attitudes, considering both robust and optimistic outcomes, can lead to better control over the balance between safety and improving outcomes. By generalizing robust decisions, tools like AAGS can be used beyond safety critical problems. Future work should focus on formal guarantees for problems with optimistic ambiguity attitudes. Furthermore, building on the work of Marinacci [22], incorporation of risk attitudes would let users independently control how agents consider both risk and ambiguity.

### v3

@@ Abstract
Due to the complexity of many decision making problems, tree search algorithms often have inadequate information to produce accurate transition models. This results in ambiguities (uncertainties for which there are multiple plausible models). Faced with ambiguities, robust methods have been used to produce safe solutions—often by maximizing the lower bound over the set of plausible transition models. However, they often overlook how much the representation of uncertainty can impact how a decision is made. This work introduces the Ambiguity Attitude Graph Search (AAGS), advocating for more comprehensive representations of ambiguities in decision making. Additionally, AAGS allows users to adjust their ambiguity attitude (or preference), promoting exploration and improving users’ ability to control how an agent should respond when faced with a set of plausible alternatives. Simulation in a dynamic sailing environment shows how environments with high entropy transition models can lead robust methods to fail. Results further demonstrate how adjusting ambiguity attitudes better fulfills objectives while mitigating this failure mode of robust approaches. Because this approach is a generalization of the robust framework, these results further demonstrate how algorithms focused on ambiguity have applicability beyond safety-critical systems.
@@ I Introduction
As autonomous agents navigate more complex environments, they are confronted with increasing model uncertainty. This is especially important with Monte Carlo tree search techniques, such as upper confidence trees (UCT) [1]. Tree search algorithms sample a blackbox simulator to produce transition models for each state encountered. However, many of these algorithms will be inaccurate due to limitations on sampling and the tree structure of the graph [2]. Consider a sailing robot, many diverse encounters and long operational periods are required before enough data is present to model commonly occurring conditions at sea. This leaves robots with too little information to reliably predict outcomes—that is to say the uncertainty is non-measurable [3] or ambiguous. Conditioned on some prior information, these ambiguities can be thought of as equally valid models of a system. For example, when sailing, we cannot precisely measure the conditions beyond the immediate area. Without this distinguishing information, we cannot predict precisely where the boat will end up prior to acting.
@@
Failure to consider these ambiguities in the context of decision making can distort the perception of uncertainty and leaves robots oblivious to potential fault conditions.
@@
Some more recent works consider ambiguity in a limited sense. These approaches generally fit under the auspices of robust decision making. Robust decision making tools construct a set of models and seek to maximize the lower bound on a state’s value [7, 8]. Representations of ambiguity in this context are limited to the use of confidence intervals [2, 9] or chance constraints on undesirable outcomes [10]. These representations neglect other information conveyed by the transition model. For example, by considering the distribution of outcomes more explicitly, we can arrive at more informed decisions [11]; coupled with representations of ambiguity we can better assess the knowledge carried by a decision maker. Additionally, the conservatism of robust methods limits their use to safety critical systems. In many other applications (e.g., service robots), giving decision makers the option to trade-off between robust policies and optimistic policies when faced with ambiguity can lead to more desirable outcomes (Fig. 1).
@@ II Ambiguity MDP
Markov decision processes (MDP) have seen broad use formulating decision problems [14] and have a substantial base of literature from which to build tools. The key issue with MDPs is the assumption of one true transition model. For methods which produce approximate solutions, this model is hidden. Depending on the approach taken, then, multiple transition models may be inferred, implying significant ambiguity when data is limited. This work builds on the concept ambiguous Markov decision processes (AMDP) [15, 16], which frame the problem with a set of transition models. Define the AMDP as $\mathcal{M}\coloneqq\langle\mathcal{S},\mathcal{A},\mathcal{T},R,\gamma\rangle$. Let $s\in\mathcal{S}$ be a state in the set of states and $a\in\mathcal{A}$ be an action in the action space. The reward $R(s,a,s^{\prime})$ defines the optimization criterion with respect to a specified state transition. The variable $\gamma$ acts a temporal discount factor. Lastly and central to the AMDP formulation, $T(s^{\prime}|s,a)\in\mathcal{T}$ is a model among the set valid of transition models. In this work the set of transition models is represented as a special class of random sets—referred to as belief functions (different from beliefs in POMDP literature [^1] Beliefs are defined specifically with respect to state uncertainty. Belief functions refer to mass functions whose elements may support more than one outcome.). As opposed to previous AMDP formulations, use of belief functions are better suited to discrete distributions than cloud models as used by Saghafian [15]. Additionally, these belief functions generalize Bayesian mass functions, allocating mass to multiple hypotheses [4]. Thus they permit more comprehensive expressions of ambiguity than those derived from Bayesian models (e.g., the AMDP of Baürle and Rieder [16]).
@@
These bounds on probability are demonstrated in Figure 2.a where each mass is known to ±$\pm\epsilon/2$. Figure 2.b captures the distribution as a belief function where mass can now support multiple hypotheses.
@@ III Methods
The goal of this work is to motivate the broader study of ambiguity in decision making. The proposed ambiguity attitude graph search algorithm builds on this by better representing ambiguity present in decision making, and generalizing robust decision making to include ambiguity attitudes. As with many sampling based methods [1, 2, 14], agents supply an $(s,a)$-pair to a simulator. This simulator returns observations in the form of a state transition and reward $\omega\coloneqq(s^{\prime},r)$. Using these observations, the agent constructs an estimate of the transition model for each $(s,a)$-pair. However, multiple underlying transition models could have produced these observations. Lacking additional information, there is ambiguity and the model estimate should instead be a set of models. Whereas prior methods, such as graph based optimistic planning (GBOP) [2], assume ambiguity is adequately described by confidence intervals, we contribute a novel way to generate belief functions from these intervals to more comprehensively represent the distribution of mass. While robust methods assume the lower bound should be used to arbitrate models, AAGS uses ambiguity attitudes to balance robustness and optimism.
@@ III-A Computing Belief Functions
Approximate decision making algorithms generally define their optimality either with respect to the budget or accuracy of estimated models [9]. For fixed confidence approaches, a confidence interval over some value is specified to accuracy $\epsilon/2$ given confidence $\delta$. The idea is that a quantity $z$ is approximated with $\hat{z}$ such that $P(|z-\hat{z}|<\epsilon/2)\geq 1-\delta$. Whereas these bounds are generally applied to the reward or value estimates, AAGS bounds the mass terms to evaluate the distribution more directly. Focusing on the entire distribution in this way has been shown to be more expressive and provide a better understanding of the model [11].
When there are limited samples, it is unlikely the constraints for a fixed confidence approach will be satisfied. We therefore develop a “floating confidence” description of the set of transition models for this case. Thus users can understand the confidence with which a distribution of desired accuracy may be interpreted. A belief function is then evaluated using linear optimization. Using the confidence, the inferred set of models is discounted against the set of all models to arrive at a distribution expressing the ambiguity.
Begin by assuming the transition model is represented by a multinomial distribution, known to $(\epsilon,\delta)$. Our approach is agnostic to how these are obtained (for clarity, implementation details are presented in the Appendix). We wish to redistribute the $\epsilon$ mass from each outcome to ambiguous mass terms, generating a set of transition models. Define $\Theta$ as the set of outcomes $(s^{\prime},r)$ from our state transition, and let $2^{\Theta}$ be the power set of all such outcomes. Let each proposition of size one be $\omega_{i}\in\Theta$ and propositions of size two or more be $\theta_{j}\in\{2^{\Theta}\setminus\Theta\}$. Let $p_{w}$ be the sampled probability of some $\omega$. The accuracy constrains belief functions such that the lower and upper bounds are, respectively, $Bel(\omega)=p_{\omega}-\epsilon/2$ and $Pl(\omega)=p_{\omega}+\epsilon/2$, $\forall\omega\in\Theta$. Because these are probabilities, they are clipped to 0 and 1, as necessary.
@@ III-B Ambiguity Attitude Graph Search
The AAGS algorithm is a graph based planner; when a state is reached multiple times, rather than creating a new node, the existing node is added as a child of the current node. In doing so, the graph using samples more efficiently than tree methods—as the decision maker can leverage more prior knowledge [2]. Assuming the simulator is much more expensive than the algorithm itself, the cost of any added computation are far outweighed by the reduced sample complexity. Cyclic graphs can occur during search and there are no constraints on connectivity.
@@ IV Results
We compare AAGS against two algorithms. First, UCT [1] serves as a baseline. Because UCT does not consider sets of models in its value estimates, it is ambiguity neutral. UCT was tuned to have a rollout horizon of 25 time steps and exploration parameter $c=8$. Next, GBOP [2] was used to compare against robust planning. Both AAGS and GBOP were given a horizon of 50 time steps. All algorithms were allocated a budget of 500 samples from the generative model.
@@ IV-A Expression of Ambiguity
First, let us draw attention to the parameter $c$ used by UCT. This exploration parameter affects its search, not the representation of value during decision making. Thus, at leaf nodes there is no concept of uncertainty in the backpropagated reward. Robust algorithms have begun to consider this uncertainty, however, they assume the lower bound is most relevant to the decision. This motivates the use of ambiguity attitudes to reflect how willing the agent should be to accept unknown gambles. One impact, is that users can control the extent to which the agent explores its environment. This is demonstrated in Figure 1 where we compare the role of planning horizons with ambiguity attitude for AAGS. In some cases long planning horizons ($h=100$) are necessary to for a robust attitude ($\alpha=0$) to identify high value targets. In which case, they move directly towards the target. Given a short planning horizon and an increasingly optimistic attitude ($h=10;\;\alpha\rightarrow 1$), however, agents increasingly seek alternatives. They move towards ambiguous regions to learn about their environment. This results in more exploratory behavior despite shorter search horizons.
@@ IV-B Sailing
We first look to the role of ambiguity in avoiding unsafe outcomes. Interestingly GBOP, a robust algorithm struggles with this (Fig. 4.a). We postulate this is because the fixed confidence models employed are relatively uninformative in environments with high entropy transitions such as this. The UCT and robust ($\alpha=0$) AAGS algorithms, however, almost completely avoid such negative outcomes. Though this number does trend up slightly with increases in initial distance (from the goal), as agents are more likely to start in the boundary. Furthermore, we see that by adjusting the ambiguity attitude of AAGS, we can directly control the willingness of an agent to gamble on low quality outcomes (Fig. 4.b). This is further reflected by the rewards (Fig. 5). For moderate attitudes, an agent can effectively balance reward seeking with chances of hitting the reefs. This can lead to improvements over UCT by $\sim$10%. Conversely, a robust attitude may significantly underperform. Similarly, an optimistic one may gamble too much and degrade outcomes.
Figure 5: (a) Reward for UCT and robust solutions (AAGS at $\alpha=0$ and GBOP) and (b) reward achieved by AAGS by varying $\alpha$. Focusing on raising the lower bound, robust attitudes achieve lower reward. Conversely, slightly positive attitudes can outperform ambiguity neutral methods (UCT). When attitudes are too optimistic ($\alpha=1$), unnecessary gambles harm performance.
@@ V Conclusions
The increasing complexity of tasks we assign to robots entails ambiguity on the part of the decision maker. This work explores how a decision maker infers multiple models. While existing methods do this in varying capacities, this work aims to demonstrate how more explicit representations of ambiguity contribute to understanding uncertain decisions. In doing so, this lays the groundwork for novel multiple model problems. Additionally, this work demonstrates how ambiguity attitudes, considering both robust and optimistic outcomes, can lead to better control over the balance between safety and improving outcomes. By generalizing robust decisions, tools like AAGS can be used beyond safety critical problems. Future work should focus on formal guarantees for problems with optimistic ambiguity attitudes. Furthermore, building on the work of Marinacci [22], could lead to methods which independently control how agents consider both risk and ambiguity.